Instructor Ramos is concerned about the amount of time teachers spend each week doing


schoolwork at home. A simple random sample of 56 teachers had a mean of 8.0 hours per week


working at home after school. Construct and interpret a 95% confidence interval for the mean


number of hours per week a teacher spends working at home. Assume that the population


standard deviation is 1.5 hours per week.



1
Expert's answer
2021-07-09T12:56:23-0400

"n=56 \\

\mu=8.0 \\

\sigma=1.5"

Two-sided confidence interval:

"CI = (\mu - \frac{Z_c \times \sigma}{\sqrt{n}}, \mu + \frac{Z_c \times \sigma}{\sqrt{n}}) \\

CI = (8.0 - \frac{1.96 \times 1.5}{\sqrt{56}}, 8.0 + \frac{1.96 \times 1.5}{\sqrt{56}}) \\

CI = (8.0 - 0.39, 8.0 + 0.39) \\

CI = (7.61, 8.39)"